OpenStudy (anonymous):

Add. 5 1/8 + 5/6 + 5/12 A.6 3/4 B.6 3/8 C.6 3/12 D.6 3/24

OpenStudy (anonymous):

Medals will be given to answers

OpenStudy (destiny_forever):

D

OpenStudy (anonymous):

Can you explain why It's d?

OpenStudy (lolacole12):

i thought it was b

OpenStudy (lolacole12):

becuase 518+56+512 =638(Decimal: 6.375)

OpenStudy (lolacole12):

6 3/8

OpenStudy (anonymous):

It could be B I'm still picking my answers. It's either D or B I might go with B but I need more opinions.

OpenStudy (lolacole12):

i thinks thats right b

OpenStudy (anonymous):

I came up with none of these... These are fractions why are you adding them together

OpenStudy (anonymous):

Fractions can be added together. Search up some results. I know that for a fact but I can't answer this because I'm confused.

OpenStudy (lolacole12):

yes im right look http://www.mathpapa.com/algebra-calculator.html 5 1/8 + 5/6 + 5/12

OpenStudy (lolacole12):

and lemme see @Deadly_Roses help

OpenStudy (destiny_forever):

im sorry its B.5 and 1/8 + 5/6 + 5/12 is the same as saying 41/8 + 5/6 + 5/12. Now we need to find a common denominator. The least common multiple of the 8, 6, and 12 is 24. So we after multiplying 41/8 with 3, 5/6 by 4, and 5/12 with 2 to get a common denominator of 24 . We get 123/24 + 20/24 + 10/24. Now we add it all together and get 153/24. Now simplify. 153/24= 6 and 9/24. But we can still simplify it further 6 and 9/24 = 6 and 3/8. You can also leave the 5 just as it is and multiply 1/8 by 3 to get the common denominator of 24. So it would be 5 and 3/24 + 20/ 24 + 10/ 24. When you add the numerators together you get 5 and 33/24. And you simplify it to 6 and 3/8.

OpenStudy (lolacole12):

thank you des see im right

OpenStudy (anonymous):

Thanks. Is there a way to give you both medals?

OpenStudy (lolacole12):

i gave her medal and u give me one

OpenStudy (anonymous):

It looks like I have to answer 3 more questions. Do I type them here or make another question?

OpenStudy (lolacole12):

yep

OpenStudy (destiny_forever):

your welcome. Have a great day:)

OpenStudy (lolacole12):

wats your ques